How do we write 1/81 as a positive integer

1
Expert's answer
2021-08-09T14:44:26-0400
"\dfrac{1}{81}\approx0.01234568"

If you insist on converting "\dfrac{1}{81}" to a whole number, the best we can do is round the decimal number up or down to the nearest whole number. "\dfrac{1}{81}" rounded up to the nearest whole number is 1 and "\dfrac{1}{81}" rounded down to the nearest whole number is 0.


Solving the equation

Multiply the numerator of the inverted fraction by the number you wish the original fraction to become, if you wish the number to be an integer other than 1.

For example,

"81(\dfrac{1}{81}x + 5 = 9) = 81(\dfrac{1}{81}x ) + 81(5) = 81(9)."

This resolves to "x+405=729." Notice that the fraction "\dfrac{1}{81}" is now the integer "1," making the equation easier to solve.


"x=324"